### v2

# DySurv: Dynamic Deep Learning Model for Survival Analysis with Conditional Variational Inference

Abstract
Survival analysis focuses on estimating time-to-event distributions which can help in dynamic risk prediction in healthcare. Extending beyond the classical Cox model, deep learning techniques have been developed which moved away from the constraining assumptions of proportional hazards. Traditional statistical models often only include static information where, in this work, we propose a novel conditional variational autoencoder-based method called DySurv, which uses a combination of static and time-series measurements from patient electronic health records to estimate the risk of death dynamically. DySurv has been tested on several time-to-event benchmarks where it outperforms existing methods, including deep learning methods, and we evaluate it on real-world intensive care unit data from MIMIC-IV and eICU. The predictive capacity of DySurv is consistent and the survival estimates remain disentangled across different datasets supporting the idea that dynamic deep learning models based on conditional variational inference in multi-task cases can be robust models for survival analysis.
Index Terms:
deep learning, healthcare, personalized medicine, prognostication, survival analysis, variational autoencoders
I Introduction
Survival analysis refers to statistical approaches to estimating distributions of event times or times it takes for an event to happen as well as rates of survival over time while accounting for censoring. The events in question can be machine failures in industry or the occurrence of specific diseases and death [1]. In clinical practice, survival analysis can play a key role and provide valuable information to predict patient outcomes and guide treatment decisions [2]. While most traditional applications of survival analysis occur in epidemiology at the population level, with the rise of deep learning techniques, personalised estimation of survival times for individual patients has become possible [3]. However, the limitations of standard statistical models such as the Cox proportional hazards model include lack of complexity, constraining assumptions about the behaviour or proportions of covariate effects over time, and the reliance on only using static covariates [4]. While there are extensions of Cox to time-varying covariates, these still suffer from the proportionality assumption (often violated in practice) and are often outperformed by simple deep learning models [5]. In settings like healthcare and with the occurrence of electronic health records, high-dimensional longitudinal information on the patient’s state can be an informative source for the prediction of clinical risk of mortality and other events.
To incorporate longitudinal, high-dimensional, and potentially multimodal data and to move away from constraining assumptions of simple statistical models, deep learning methods have been proposed in survival analysis. These include extensions to the Cox model which estimate the parameters using a deep learning model that thus remain restrained with the proportionality assumption as in the case of DeepSurv [6]. Other deep learning models include DeepHit and its extension to longitudinal data, Dynamic-DeepHit, which implements a custom loss and a simple recurrent neural network to avoid making any parametric assumptions on the survival distribution [7, 8]. Simple extensions of recurrent neural networks for estimating survival distributions have also been proposed [9, 10]. More recently, the autoencoder structure has been proposed to learn from a combination of data modalities by finding a lower-dimensional latent representation of the data. ConcatAE uses a simple autoencoder on multi-omics data to find a hidden representation then used in a multi-class classification task to predict discrete survival times [11]. This treats the survival analysis task as one of simple classification and does not aim to estimate the survival distribution at all. On the other hand, the variational autoencoder present in VAECox, directly estimates the survival distribution using the Cox loss function while taking advantage of the variational inference over traditional autoencoders [12]. Due to the reliance on the Cox loss, VAECox suffers from the often-violated proportionality assumption meaning its results in real-world datasets could be limited.
Our work addresses these limitations by using a cumulative incidence risk estimation loss function based on the negative log-likelihood which requires making no parametric or proportionality assumptions on the survival distribution and hazard risks. We also extend the variational setting to conditional VAE which we find improves predictive performance. Our model, DySurv, is compatible with both static and longitudinal time-series data enabling more comprehensive learning from patient electronic health records. We validate our approach both in static and time-varying settings using six benchmark datasets as well as two large public open-access ICU datasets. The critical care setting provides high-frequency longitudinal measurements and short-term health outcomes in which dynamic risk stratification of survival analysis could provide additional insights [13]. We aim to show that conditional variational inference and autoencoder reconstruction tasks can improve learning from complex time-series data by extracting latent features to optimise the survival task. Since DySurv is not only multimodal in input but also in output with both reconstruction and survival tasks included, conditional variational autoencoders (CVAE) provide better predictive performance [14]. By mapping high-dimensional covariates into a lower-dimensional space, CVAEs can effectively capture the underlying structure and identify hidden factors that significantly impact survival outcomes to predict mortality as well as time-to-death. This capability allows deeper insights into dynamic mortality risk prediction in healthcare.
II Materials and Methods
II-A Data
Standard benchmark datasets contain only static features but here we implement survival analysis on the ICU datasets from MIMIC-IV and eICU which contain both static and time-series data. To show performance across different datasets and with different sizes, we will succinctly introduce these datasets. Across all datasets, the event in question is death. The datasets were split into 60% training, and 20% each for validation and testing. Quantile transformations have been applied for standardisation and fit only on the training dataset. Please see the appendix for a full list of features for each dataset.
II-A1 SUPPORT
The Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments contains data from five care academic centres in the United States for patient survival in the following six months [15]. The result of the study was a prognostic model to estimate survival for seriously ill hospitalised patients. The dataset consists of 8,873 samples and 14 static features.
II-A2 METABRIC
The Molecular Taxonomy of Breast Cancer International Consortium contains genetic and clinical data from breast cancer patients with 1,904 samples and 9 static features [16].
II-A3 GBSG
The Rotterdam & German Breast Cancer Study Group contains treatment and clinical data on 2,232 breast cancer patients with 6 static features [17].
II-A4 NWTCO
The National Wilm’s Tumor dataset contains staging and clinical data on 4,028 Wilms’ tumour patients with 6 static features [18].
II-A5 sac3
The simulated dataset contains discrete event times with 44 static features and 100,000 samples [19].
II-A6 sac_admin5
The simulated dataset contains discrete event times with 5 static features and 50,000 samples [20].
II-A7 MIMIC IV and eICU
We conduct experiments on the de-identified real-world ICU dataset Medical Information Mart for Intensive Care (MIMIC-IV v. 2.0, July 2022) including discharge information for more than 15,000 additional ICU patients compared to the previous release [21]. The dataset contains data from Beth Israel Deaconess Medical Center collected between 2008 and 2019. The dataset contains 71,935 samples of ICU stays with 33 static features (categorical features were one-hot encoded) and 65 time-varying features. Furthermore, we tested survival analysis models in the eICU Collaboration Research Database [pollard2018eicu]. The eICU database was processed using postgreSQL and the pandas package. eICU is a multi-centre ICU database with over 200,859 patient unit encounters for 139,367 unique patients admitted between 2014 and 2015 to one of 335 ICUs at 208 hospitals located throughout the United States. The database is de-identified and includes vital sign measurements, demographic data, and diagnosis information. Static variables include age, sex, admission unit, and others that did not have missingness. For the time-series variables, we use forward filling as clinicians in practice would only consider the last recorded measurement. If the first set of measurements is missing for some time-varying features, instead of dropping those features or patients, we backward fill from the closest measurement in the future. The processing of MIMIC-IV follows from previously published work on this dataset but is adapted for the survival scenario with the duration of stay in the ICU or the maximum time horizon for the event times defined as 10 days and time-series features taken in 72-hour timesteps [22]. Each node of the neural network model then corresponds to a 1-day risk prediction in the ICU as the maximum time horizon is 10 days. Censoring, as always, is defined as loss-to-follow-up or discharge.
II-B DySurv
In survival analysis, the main underlying goal is the estimation of the survival function which represents the probability that no event occurs until a time t and can be written as
[…]
where $f(t)$ is the probability density function of event time and
$f(t)=\lim_{\Delta t\rightarrow 0}\frac{\operatorname{P}(t<T\leq t+\Delta t)}{\Delta t}$
and $F(t)$ corresponds to the cumulative risk or incidence function $F(t)=\operatorname{P}(T\leq t)$. $T$ represents the time of the event, $\operatorname{P}$ is the probability, and $t$ is the specitic timestamp for risk estimation. The key to training a deep learning model is to learn an estimate of the cumulative incidence or risk function $\hat{F}(t)$ as the joint distribution of the event time and outcome label given the observations. In general, the time-to-event values can be left to be continuous depending on the model being considered but we discretise the time set into 10 equally spaced time periods for each dataset in the fashion of DeepHit and Dynamic-DeepHit [8]. As we discretise the time into intervals, we can estimate this event probability across arbitrary periods and remain faithful to the original survival analysis formulation rather than resorting to chained binary classification [23]. Furthermore, measurements are often right-censored, meaning that patients can leave the study or be lost to follow-up and they therefore do not experience an event at all. A common assumption is that this censoring is not important and independent of the outcome of the study itself [24]. Once we have an estimate of the risk, we can then simply obtain the survival function and curves by taking its negation from one as shown in equation (3).
Data for survival analysis contain three main sets of variables, the first is the feature set which can consist of static or time-series features (the latter having measurements at potentially different sampling frequencies), the time-to-event for the events in question or censoring, respectively, and the outcome label for an event [25]. Thus, the dataset can be represented as
$\mathcal{D}=\left\{\left(\mathcal{X}^{i},\mathcal{t}^{i},y^{i}\right)\right\}_{i=1}^{N}$ (2)
with $\mathcal{X}^{i}$ representing the feature matrix of patient $i$, $\mathcal{t}$ being the time-to-event ie. the minimum of the event and censoring time, $y$ being the label for the outcome, and $N$ samples included. A feature matrix (in the time-series case) for patient $i$, $\mathcal{X}^{i}$, can be seen as
$\mathcal{X}^{i}=\left\{\boldsymbol{x}_{1}^{i},\boldsymbol{x}_{2}^{i},\ldots,\boldsymbol{x}_{j}^{i}\right\}$ (3)
where $j$ is the length of the time-series for $1\leq j<J^{i}$ where $J$ is the maximum time step with timestamps of measurements $\left[t_{1}^{i},t_{2}^{i},\ldots,t_{j}^{i}\right]$. $\boldsymbol{x}_{j}^{i}$ contains $M$ set of features $\left[x_{j,1}^{i},x_{j,2}^{i},\ldots,x_{j,M}^{i}\right]^{\top}$ for timestamp $t_{j}^{i}$.
We undertake estimation of the underlying cumulative risk by mainly optimizing the negative logarithmic likelihood of the joint distribution of the event time and outcome with right-censoring. For those patients who have suffered the event, we capture both the outcome and the time at which it occurs. For censored patients, we capture the censoring time conditioned on the measurements recorded prior to the censoring. If we assume $\hat{a}_{t}=P\left(T=t\mid\mathcal{X}\right)$ (the output of the last layer node of the neural network module) represents the estimated probability of experiencing an event at time $t$, then the loss can be represented as
$\mathcal{L}_{1}=-$ $\sum_{i=1}^{N}\left[\mathbb{1}\left(y^{i}\neq\varnothing\right)\cdot\log\left(\frac{\hat{a}_{t^{i}}^{i}}{1-\sum_{y\neq\varnothing}\sum_{n\leq t_{Ji}^{i}}\hat{a}_{n}^{i}}\right)\right.$ (4)
$\left.+\mathbb{1}\left(y^{i}=\varnothing\right)\cdot\log\left(1-\sum_{y\neq\varnothing}\hat{F}\left(t^{i}\mid\mathcal{X}^{i}\right)\right)\right]$
where $\mathbb{1}$ is the indicator function. The first term represents optimising for uncensored patients and the second for those censored at the last noted time ie. alive. The estimated cumulative incidence risk, $\hat{F}$, is the probability that the patient dies at time $t$ or before, conditioned on all previous longitudinal measurements. By optimising for this loss, we estimate the actual risk distribution for each patient and a prediction can be made for arbitrary times across event times as outputs of the last layer.
As data are complex (i.e., static and time-varying), we seek to learn a latent representation using a conditional variational autoencoder that would improve learning for the task of survival analysis. There are two loss terms, namely the reconstruction (mean squared error) and the Kullback–Leibler (KL) divergence in the variational autoencoder. Estimating the distribution of the underlying latent factors relies on minimising the KL divergence between an approximation of the true posterior and the true distribution both of which are assumed to be multivariate Gaussians. As such, learning these distributions means optimising for their parameters, the mean and standard deviation. $z$ is the sampled latent vector from the probabilistic encoder for the learned Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. For training, however, since sampling is a stochastic process, we use the reparameterization trick to backpropagate the gradient and represent the latent vector as the sum of a deterministic variable and an auxiliary independent random variable $\varepsilon$ [26]
$\varepsilon\sim N(0,1)\quad z=\mu+\varepsilon*\sigma\quad\rightarrow\quad z\sim N(\mu,\sigma)$ (5)
The ability of the decoder to successfully reconstruct the input is captured with a simple mean squared error term between the reconstruction of the input and the input itself. Thus, the loss for variational inference can be seen as
$\mathcal{L}_{2}=L_{vae}(E,D)=\left\|\mathcal{X}-\mathcal{X}_{\text{recon }}\right\|_{2}$ (6)
$+\frac{1}{2}\sum_{i=1}^{z_{\text{dim }}}[\left(\mu_{i}^{2}+\sigma_{i}^{2}\right)-1-\log\left(\sigma_{i}^{2}\right))]$
The latent vector $z$ is then used as input to a neural network module in optimising the survival task. During training, the decoder uses the latent vector and the condition vector (survival labels or times of death) as input which helps the latent space to capture other information instead of trying to better reconstruct the input. At test time, the decoder is not used anymore, and the latent space is used for prediction in the survival task. The total loss can then be presented as
$\mathcal{L}=\alpha\mathcal{L}_{1}+(1-\alpha)\mathcal{L}_{2}$ (7)
where $\alpha$ is the balancing coefficient between the two losses ($\mathcal{L}_{1}$ and $\mathcal{L}_{2}$) and $0\leq\alpha\leq 1$. $\alpha$ is considered as a hyperparameter that is optimised during training according to multi-objective optimization principles.
Figure 5 in the Appendix shows the structure of the proposed model. DySurv leverages the dynamic nature of deep learning time-series models combined with conditional variational autoencoders for multi-task learning of survival analysis risk prediction extending beyond the classical fixed binary event prediction from traditional machine learning models. Figure 1 shows the framework employed here built around DySurv to support robust risk estimation for clinical events such as death. Using a simple autoencoder has been shown to lead to overfitting and imbalanced learning of the reconstruction task that could harm learning the survival task whereas a VAE’s objective function is based on the reconstruction loss from a randomly sampled vector allowing for more robustness [12]. We concatenate the feature vectors from the static and time-series features together before feeding them into an encoder equipped with a Long-Short Term Memory (LSTM) cell.
The time-series components are compressed into a latent representation that is then used, as is, as input to the estimation of the negative log-likelihood loss function. A detailed description of the architecture can be found in the Appendix. Once the latent vector is sampled from the Gaussian distribution defined by these parameters, the lower-dimensional latent factors are used as input for an MLP module to optimize the survival task. The hyperparameters optimised in the network through grid search using the training and validation set included learning rate, batch size, $\alpha$, and dropout proportion. To minimise overfitting, we employ early stopping techniques.
Figure 1: (a) Description of the proposed DySurv framework using longitudinal EHR data for dynamic risk prediction instead of fixed-point event classification. The patient stay consists of measurements until $t_{j}$ as the last measurement recorded in the observation window. The orange marker states how much of the longitudinal data, i.e., how many timestamps are used in the model learning process. The red marker indicates how many timestamps into the future the model estimates the risk of death. In the first instance of classical machine learning, only the most recent timestamp measurement is used to predict the risk of death at a fixed time point in the future at a prediction window distance (except for Gaussian processes). For deep learning classification, by using LSTMs, we can learn from the entire patient longitudinal measurement but only estimate the risk of death at a fixed timepoint, hence the observation window is fully orange, but only one of the prediction window timestamps is red. DySurv uses all of the available longitudinal data to predict risk dynamically, ie. at all reasonable times into the feature.
We compare the performance of our model to a collection of survival analysis models summarised in Table I and described in detail in the Appendix.
TABLE I: Survival analysis methods investigated
Method Time scale Reference
PMF discrete-time [19]
MTLR discrete-time [27]
BCESurv discrete-time [20]
DeepHit discrete-time [7]
Logistic Hazard discrete-time [28]
CoxTime continuous-time [5]
CoxCC continuous-time [5]
DeepSurv continuous-time [6]
PCHazard continuous-time [19]
II-C Metrics
In this section, we will switch the notation of samples from superscript to subscript, hence $\mathbf{x}^{i}$ is now $\mathbf{x}_{i}$ for sample $i$. Since we are no longer making single risk predictions at specific times and estimating distributions of event times for censored samples, different evaluation metrics must apply than those used in classic machine learning classification and prediction settings. The most common metric for evaluating survival analysis models is the concordance index $C_{ind}$, which estimates the probability that, for a random pair of samples, the predicted survival times (risk probabilities) of the two samples have the same ordering as their true survival times [5]. This explanation works perfectly for settings of proportional hazards where the ordering does not change over time but for our purposes, we will not be limited by such assumptions. Hence, we will rely on using the time-dependent extension $C_{ind}^{\mathrm{td}}$ with some modifications accounting for predictions independent of feature observations having a concordance of 0.5. The metric can be represented as
$C_{ind}^{\mathrm{td}}=\mathrm{P}\left\{\hat{S}\left(T_{i}\mid\mathbf{x}_{i}\right)<\hat{S}\left(T_{i}\mid\mathbf{x}_{j}\right)\mid T_{i}<T_{j},D_{i}=1\right\}$ (8)
where $\hat{S}$ indicates the estimated survival probabilities are used and $D_{i}=1$ that only those who experienced the event are considered in this metric. A noted limitation of this metric is its obvious bias and dependence on the censoring distribution as only non-censored samples are considered making it affected by the length of stay and the censoring proportion that increases over the length of stay. To this end, we decided to use additional metrics for more holistic evaluation especially as previously proposed models like DeepHit were found to have lackluster results in real-world datasets when evaluated using other metrics besides concordance. We also evaluate our model using the Integrated Brier Score or IBS. The Brier Score is similar to the mean squared error as it represents the average squared distances between the predicted and the true survival probability (approximated with step functions with jumps at the event times) and is always a number between 0 and 1, with 0 being the best possible value [29]. The expectation of the Brier Score contains the mean squared error as one of its additive terms, so minimising one is minimising the other [20]. Since we need to know the event times for calculating IBS and we do not have access to all the samples’ event times in right-censoring, an adjusted metric called the inverse probability of censoring weights Brier Score (IPCW) is used instead to approximate the times by weighting the scores of the observed event times by the inverse probability of censoring. The equation used is thus
$\operatorname{BS}_{\operatorname{IPCW}}(t)=\frac{1}{n}\sum_{i=1}^{n}[\frac{\hat{S}_{i}(t)^{2}1\left\{T_{i}\leq t,D_{i}=1\right\}}{\hat{G}_{i}\left(T_{i}-\right)}+$ (9)
$\frac{\left[1-\hat{S}_{i}(t)\right]^{2}1\left\{T_{i}>t\right\}}{\hat{G}_{i}(t)}]$
where $\hat{G}_{i}(t)=\mathrm{P}\left(C_{i}^{*}>t\right)>0$ is the Kaplan-Meier estimate of the censoring distribution for sample $i$ and $C_{i}^{*}$ is the censoring time. The expected value of this metric is the same as that for the uncensored Brier Score. As one notices, this metric is evaluated at specific times, whereas the Integrated Brier Score or IBS provides a general evaluation of model performance at all times.
[…]
A limitation of this metric, however, is the biased assumption of the censoring distribution being the same across samples thereby disregarding covariate effects. This can be addressed by using an administrative extension of the metric that requires access to all the censoring times but a discussion of this in greater detail can be perused here [20]. Lastly, we introduce the IPCW (negative) binomial log-likelihood or NBLL from classic binary classification which measures both discrimination and calibration of the estimates and uses its integrated extension for all times INBLL
$\operatorname{BLL}(t)=\frac{1}{N}\sum_{i=1}^{N}[\frac{\log\left[1-\hat{S}\left(t\mid\mathbf{x}_{i}\right)\right]1\left\{T_{i}\leq t,D_{i}=1\right\}}{\hat{G}\left(T_{i}\right)}+$ (11)
$\frac{\log\left[\hat{S}\left(t\mid\mathbf{x}_{i}\right)\right]1\left\{T_{i}>t\right\}}{\hat{G}(t)}]$
[…]
For both of the last metrics, we approximate the integrals by numerical integration (for 100 timesteps as based on previous literature), and the time span is the duration of the test set as these metrics are only evaluated on the test set [5].
III Results
To holistically evaluate DySurv, we present a set of experiments and comparisons with other benchmark survival analysis models across multiple datasets. We present not only the discriminative performance of the model as measured by concordance but also its calibration as measured by IBS and IBLL. We evaluated the model on a case study example of patients in the ICU by including the real-world MIMIC-IV and eICU electronic health record datasets. The results consist of two major experiments, one is the ability of the model to successfully learn from static data which is present in all the datasets, and the other to learn from a combination of static and time-varying data such as in MIMIC-IV and eICU. For these purposes, Dynamic-DeepHit is the only relevant comparison as other survival analysis models deal only with static data. Tables II, III, and IV show these results across datasets on held-out test sets for all models included. We implemented these methods in PyTorch (PyCox) v. 1.10.1 using a MacBook Air M1 2021 laptop with data processing completed using pandas and SQL.
TABLE II: Test results on different static-only datasets for survival analysis models and DySurv as evaluated by three different metrics introduced in Materials and Methods. For concordance, higher is better, and for the other two metrics, lower is better. The best results are in bold. All of the results are an average of five random seeds.
$C_{ind}^{\mathrm{td}}$ IBS IBLL $C_{ind}^{\mathrm{td}}$ IBS IBLL
SUPPORT METABRIC
PMF 57.9 0.195 0.574 PMF 63.8 0.168 0.497
MTLR 55.3 0.205 0.775 MTLR 56.8 0.172 0.527
BCESurv 55.3 0.290 2.08 BCESurv 56.8 0.138 0.477
DeepHit 57.3 0.273 0.678 DeepHit 65.5 0.123 0.415
Logistic Hazard 53.5 0.206 0.762 Logistic Hazard 59.0 0.163 0.498
CoxTime 59.5 0.193 0.565 CoxTime 65.4 0.114 0.361
CoxCC 59.7 0.192 0.563 CoxCC 65.9 0.166 0.508
DeepSurv 60.6 0.190 0.559 DeepSurv 62.4 0.176 0.541
PCHazard 55.1 0.206 0.633 PCHazard 51.4 0.160 0.547
DySurv 64.7 0.190 0.561 DySurv 64.5 0.120 0.387
GBSG NWTCO
PMF 68.5 0.179 0.528 PMF 69.7 0.122 0.389
MTLR 65.6 0.180 0.542 MTLR 66.8 0.109 0.403
BCESurv 65.6 0.156 0.481 BCESurv 69.1 0.108 0.393
DeepHit 68.1 0.174 0.514 DeepHit 71.1 0.118 0.348
Logistic Hazard 67.4 0.179 0.537 Logistic Hazard 66.5 0.108 0.396
CoxTime 68.4 0.171 0.510 CoxTime 70.7 0.110 0.343
CoxCC 59.6 0.205 0.597 CoxCC 70.3 0.110 0.373
DeepSurv 68.5 0.180 0.531 DeepSurv 68.3 0.115 0.391
PCHazard 55.8 0.182 0.574 PCHazard 60.2 0.118 0.465
DySurv 70.4 0.164 0.499 DySurv 70.3 0.111 0.347
sac3 sac_admin5
PMF 74.3 0.125 0.391 PMF 71.5 0.124 0.387
MTLR 65.0 0.124 0.539 MTLR 65.7 0.122 0.520
BCESurv 67.8 0.163 0.586 BCESurv 68.4 0.164 0.505
DeepHit 74.2 0.184 0.527 DeepHit 71.6 0.186 0.396
Logistic Hazard 72.0 0.120 0.492 Logistic Hazard 70.7 0.118 0.481
CoxTime 78.7 0.117 0.362 CoxTime 78.5 0.117 0.362
CoxCC 76.4 0.124 0.384 CoxCC 76.7 0.122 0.381
DeepSurv 76.1 0.126 0.390 DeepSurv 77.4 0.119 0.371
PCHazard 64.0 0.135 0.514 PCHazard 65.1 0.123 0.503
DySurv 80.6 0.112 0.359 DySurv 79.6 0.116 0.361
TABLE III: Test results on MIMIC-IV dataset for survival analysis models evaluated by three different metrics. For concordance, higher is better, and for the other two metrics, lower is better. All of the results are an average of five random seeds.
$C_{ind}^{\mathrm{td}}$ IBS IBLL
PMF 50.9 0.126 0.389
MTLR 52.4 0.126 0.389
BCESurv 52.2 0.157 0.473
DeepHit 54.4 0.137 0.421
Logistic Hazard 52.6 0.122 0.396
CoxTime 53.1 0.122 0.337
CoxCC 52.9 0.123 0.393
DeepSurv 54.2 0.128 0.403
PCHazard 51.0 0.122 0.378
DySurv (static) 55.7 0.111 0.360
Dynamic-DeepHit 56.0 0.143 0.376
DySurv ($+$ time-series) 57.9 0.122 0.320
TABLE IV: Test results on eICU dataset for survival analysis models evaluated by three different metrics. For concordance, higher is better, and for the other two metrics, lower is better. All of the results are an average of five random seeds.
$C_{ind}^{\mathrm{td}}$ IBS IBLL
PMF 53.4 0.115 0.323
MTLR 54.2 0.116 0.344
BCESurv 53.1 0.138 0.416
DeepHit 57.9 0.126 0.332
Logistic Hazard 54.7 0.111 0.358
CoxTime 55.7 0.112 0.310
CoxCC 55.1 0.109 0.358
DeepSurv 54.2 0.128 0.403
PCHazard 53.1 0.103 0.346
DySurv (static) 58.1 0.103 0.322
Dynamic-DeepHit 58.8 0.131 0.352
DySurv ($+$ time-series) 60.3 0.102 0.319
In survival analysis, it is also vital to show the performance of the models in creating survival curves, or estimates of the survival probabilities over time for different patients. Whereas traditional survival analysis models from statistics rely on risk sets and computing survival estimates for a population, an advantage of deep learning models is that the task of survival estimation can apply to an individual patient sample. We provide survival curves for a random group of five samples/patients (with different event times) across all datasets that show a clear separation of risk as can be seen in Figure 2. For MIMIC-IV, the data input consists of both static and time-series data and we provide survival curve results for both scenarios as well as for Dynamic-DeepHit for comparison. DySurv has a better capacity to separate independent patient risk trajectories as shown in Figure 3.
(a) Survival curve (estimate of survival probability over time) for five random samples from SUPPORT
(b) Survival curve (estimate of survival probability over time) for five random samples from METABRIC
(c) Survival curve (estimate of survival probability over time) for five random samples from GBSG
(d) Survival curve (estimate of survival probability over time) for five random samples from NWTCO
(e) Survival curve (estimate of survival probability over time) for five random samples from sac3
(f) Survival curve (estimate of survival probability over time) for five random samples from sac5
Figure 2: The set of survival curves for benchmark datasets as generated by DySurv shows clear extrapolation of risk across different samples. Since DySurv provides discrete estimates over time-intervals, the curves were generated with additional interpolation commonly used in other discrete methods in survival analysis.
(a) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using only static features with DySurv
(b) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using both static and time-series features with DySurv
(c) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using both static and time-series features with Dynamic-DeepHit
Figure 3: The set of survival curves for the MIMIC-IV ICU EHR dataset as generated by DySurv that shows clear extrapolation of risk across different patients as compared to both static, time-series, and Dynamic-DeepHit cases. The same set of patients was used in generating the survival curves across cases.
Upon deployment of a trained model, DySurv can generate risk estimates through time for each patient while using their history of observations. We do not rely on landmarking methods or a specific pre-defined time for risk prediction as the scores are simultaneously issued across the entire time interval. For data pre-processing purposes, a time horizon is selected corresponding to 24 hours with a 72-hour timespan for the LSTM in the case of MIMIC-IV and eICU since ICU risk assessments often use information over 72 hours for the next 24-hour risk prediction [30]. We also compare DySurv for 24-hour risk prediction to existing ICU survival scores in practice like APACHE IV and SOFA scores which can be seen in Figure 4.
Figure 4: Comparison of DySurv at 24-hour prediction with existing ICU survival scores using (a) AUROC and (b) Sensitivity. APACHE IV was only able to be retrieved from MIMIC-IV and SOFA score was calculated from the eICU dataset. The calculations are available in the code repository.
IV Discussion
The first set of results from Table II relate to applying DySurv only on static data from several benchmark datasets of varying sizes. We see that for the vast majority of these benchmarks, DySurv outperforms both standard statistical as well as deep learning alternatives across all metrics except for METABRIC and NWTCO on the concordance where DeepHit tends to perform slightly better. This is probably due to the implementation of the biased ranking loss mentioned earlier that aids in having better discriminative performance as measured by the concordance metric but that is not reflected as measured by the other two metrics. Similar behaviour for DeepHit has been observed in another study by [5]. We also see that the non-VAE implementation of the logistic hazard performs much worse than DySurv across all experiments, thereby strengthening the idea that adding variational inference to the logistic hazard can aid in learning the survival task. This improvement occurs despite having the additional task of reconstruction now. The predictive advantage comes from the identification of lower dimensional latent vectors used in the survival task instead of the raw features directly. Furthermore, on very large synthetic datasets, such as sac3 and sac_admin5, DySurv performs better due to having a greater amount of data to learn from. A limitation of the other benchmark datasets is their relatively small size may constrain DySurv from learning its optimal parameters to provide better survival prediction.
While there are previous models that have attempted to use autoencoders for survival analysis, such as [12] and [11], they have not explored variational inference extensively. These models also rely on optimisation of the Cox partial log-likelihood loss, hence being restricted by the proportionality assumption and they do not account for dynamic time-series or time-varying features in the input. Furthermore, work by [12] suggests that the VAE model’s learned compact latent representation directly aids in the improved performance of the Cox model. This intuition is precisely what we have also seen in our results, albeit in a larger, more flexible, and more complex scenario of time-series ICU risk prediction with direct joint distribution estimation. The concatenation autoencoder from [11] is not even compatible with static data, by far the most common modality in survival analysis, thereby limiting its relevance significantly. DySurv addresses all these limitations and provides a flexible solution to dynamic survival analysis with deep learning.
As for MIMIC-IV and eICU results, we see that even with only static data in the input, DySurv manages to outperform other survival analysis models. When time-series data is added to the input in a multi-modal fashion, it can help the model improve its performance and outperform both CoxTime and Dynamic-DeepHit across all metrics. The survival curves in Figure 3 show clearly that the static version of DySurv has the clearest separation of survival trajectories for different patients and that the addition of time-series data makes the task more complex. This is probably due to the reconstruction task being a lot more difficult now that time-series are involved and the model makes sacrifices on the survival task front. However, compared to Dynamic-DeepHit, whose survival curves in Figure 3 c) are inseparable, the time-series version of DySurv still provides a relatively reliable survival estimate. We also see that, as we would expect clinically, the survival of the patients significantly changes in the last few days in the ICU, starting to drop a few days before death. Previous work has shown that earlier times in the ICU correspond to higher survival rates. This suggests that identifying the time period when survival rates drop dramatically in the ICU can help target earlier treatment for those most at risk [31].
V Conclusion
In this paper, we present a novel dynamic risk prediction model for survival analysis based on deep learning in survival analysis paradigms. Our method builds on a combination of previous work including Dynamic-DeepHit by leveraging direct learning of the joint distribution of the first event time and the event through log-likelihood optimisation with logistic hazards. Theoretically, this approach is an alternative to the risk log-likelihood loss function of Dynamic-DeepHit itself that does not use a ranking loss for biased inflation of concordance results. Our DySurv model is capable of learning from complex EHR ICU time-series data and extracting lower-dimensional latent representations that can be useful for learning the survival task while also balancing reconstruction. As the model has been difficult to train due to loss instabilities and sensitivity to hyperparameter selection, future work can explore including a regularisation component to the loss terms. By using the underlying latent distribution, we can directly model an alternative to the survival distribution like the Weibull distribution instead of using a Gaussian intermediate.
VI Conflicts of interest
The authors declare no competing interests.
VII Code and Data sharing
The data used in the study are all publicly available. MIMIC-IV can be accessed after some initial requests and approvals from the PhysioNet website. The remaining datasets have their sources listed in the PyCox library repository. All of the code will be made available in a public GitHub repository after publication, including easily accessible notebooks to reproduce experiments from the paper.
VIII Acknowledgements
Munib Mesinovic appreciates the support of the Rhodes Trust and the EPSRC Centre for Doctoral Training in Health Data Science. T.Z. was supported by the Royal Academy of Engineering under the Research Fellowship scheme.
Appendix A Appendix
A-A Data Description
A detailed list of features used in the study from different datasets can be seen in Tables V, VI, VII, VIII and IX.
TABLE V: Features extracted from the SUPPORT survival dataset.
Feature Type Feature Type
Sex binary Age integer
Number of comorbidities integer SUPPORT Coma Score continuous
Years of Education continuous APS III Score continuous
Diabetes binary Dementia binary
3rd Day MAP continuous 3rd Day WBC continuous
3rd Day HR continuous 3rd Day RR continuous
Temperature continuous 3rd Day Bilirubin continuous
TABLE VI: Features extracted from the NWTCO survival dataset.
Feature Type Feature Type
Relapse binary Age integer
Histology Local binary Histology Institute binary
Tumor Stage categorical Study Type categorical
TABLE VII: Features extracted from the GBSG survival dataset.
Feature Type Feature Type
Hormone Treatment binary Age integer
Tumor Grade categorical Menopause Treatment binary
Progesterone (fmol) continuous Number of positive nodes categorical
Estrogen (fmol) continuous
TABLE VIII: Features extracted from the METABRIC survival dataset.
Feature Type Feature Type
Age integer Gene I binary
Gene II binary Gene III binary
Hormone Treatment binary Gene IV binary
Radiotherapy binary Chemotherapy binary
ER-positive binary
TABLE IX: Features extracted from the MIMIC-IV database. The features include demographic data collected for all patients, ICU unit-specific information like the type of unit, hospital information like regional location, time since admission, vital signs including respiratory rate and blood pressure, and biochemical measurements including blood glucose and haemoglobin.
Static Variables
Feature Type Feature Type
Sex binary Motor Response continuous
Age integer Verbal Response continuous
Height continuous Ethnicity categorical
Weight continuous Unit Type categorical
Hour of Admission integer Admission Location categorical
Time Since Admission continuous
Time-series Variables
Feature Type Feature Type
GCS - Eye continuous
Eye Response continuous GCS - Motor continuous
Braden Score continuous GCS - Verbal continuous
Strength L Arm continuous Daily Weight continuous
Strength R Arm continuous ALT continuous
Strength L Leg continuous AST continuous
Strength R Leg continuous HCO3 continuous
Insurance categorical Hct continuous
ALT continuous Alkaline Phosphatase continuous
Anion Gap continuous AST continuous
Base Excess continuous Bicarbonate continuous
Bilirubin continuous Calcium continuous
Total CO2 continuous Chloride continuous
Creatinine continuous Glucose continuous
Hematocrit continuous Hemoglobin continuous
INR(PT) continuous Lactate continuous
MCH continuous MCHC continuous
MCV continuous Magnesium continuous
PT continuous PTT continuous
Phosphate continuous Platelet Count continuous
Potassium continuous RDW continuous
Red Blood Cells continuous Sodium continuous
Urea Nitrogen continuous White Blood Cells continuous
pCO2 continuous pH continuous
pO2 continuous JH-HLM continuous
Dyspnea Assessment continuous Daily Weight continuous
Glucose continuous Heart Rate continuous
DBP continuous SBP continuous
O2 Flow continuous O2 Sat (%) continuous
Pain Level continuous Pain Level Response continuous
Phosphorous continuous Respiratory Rate continuous
Richmond-RAS Scale continuous Temperature (${}^{\circ}$) continuous
TABLE X: Features extracted from the eICU database. The features include demographic data collected for all patients, ICU unit-specific information like type and number of beds, hospital information like regional location and teaching status, vital signs including respiratory rate and blood pressure, and biochemical measurements including troponin and levels of potassium and protein in the blood.
Feature Type Feature Type
Sex binary Unit Stay Type categorical
Age integer Num Beds Category categorical
Height continuous Region categorical
Weight continuous Teaching Status binary
Ethnicity categorical Physician Speciality categorical
Unit Type categorical Unit Type categorical
Unit Admit Source categorical Mechanical Ventilation binary
Unit Visit Number categorical
Time-series Variables
Feature Type Feature Type
Base Excess continuous
-basos continuous FiO2 continuous
-eos continuous HCO3 continuous
-monos continuous Hct continuous
-polys continuous Hgb continuous
ALT continuous MCH continuous
AST continuous MCHC continuous
BUN continuous MCV continuous
O2 Sat (%) continuous MPV continuous
PT-INR continuous PT continuous
RBC continuous PTT continuous
RDW continuous WBC continuous
Alkaline ph. continuous Albumin continuous
Bedside Glucose continuous Anion Gap continuous
Calcium continuous Bicarbonate continuous
Creatinine continuous Glucose continuous
Lactate continuous Magnesium continuous
pH continuous paCO2 continuous
paO2 continuous Phosphate continuous
Platelets continuous Potassium continuous
Sodium continuous Bilirubin continuous
Protein continuous Troponin - I continuous
Urinary s. Gravity continuous mean BP continuous
SBP continuous DBP continuous
A-B Benchmark Models
To compare DySurv to other survival analysis methods, we implement a selection of the most popular and consistently cited methods in the field and provide a short description of each below. We will first introduce discrete-time methods which rely on discretising the event times into specified durations, and then follow with continuous-time methods.
A-B1 PMF
The parametrisation of the Probability Mass Function (PMF) of the event times is another way of estimation without resorting to using discrete-time risk or hazard in likelihood optimisation. We described the continuous probability density function without using discrete time boundaries. It is the foundation of other methods like DeepHit and Multi-Task Logistic Regression. It similarly resorts to optimising a negative log-likelihood loss but instead of using the cumulative risk function, it uses the approximations of the PMF and survival functions [19]. Since we can establish a direct representational relation between the risk and survival functions, the PMF loss can be seen as an alternative to our loss function. In our and others’ implementation, the PMF method is a simple Multi-Layer Perceptron (MLP) optimised for this loss and we use the same structure as we use for our survival module to be described later.
A-B2 MTLR
Multi-Task Logistic Regression provides a generalization of the binomial log-likelihood to jointly model the sequence of binary labels for each time interval risk prediction. This method similarly minimises the negative log-likelihood with the PMF and survival function terms but the network outputs are cumulatively summed in reverse to no certain advantage and, in fact, just add computational complexity and numeric instability [27, 29].
A-B3 BCESurv
This Binary Cross-Entropy for Survival is a method consisting of a set of binary classifiers that remove individuals as they are censored. The loss is the binary cross entropy of the survival estimates at a set of discrete times, with targets that are indicators of surviving each time. Each output node in the last layer corresponds to a binary classifier evaluated at that time point. As censored patients are removed, the method is biased towards those with higher event probabilities [20].
A-B4 DeepHit
The single-risk version of DeepHit is a deep learning model whose output nodes are softmaxed to jointly model the probabilities between the event times and the time durations are discretised like in our case. The model depends on optimising both the negative log-likelihood loss based on the cumulative incidence function and a ranking loss built on the intuition of the concordance. The ranking loss penalises incorrect ordering of patient pairs in which the patient that remains longer in the study should have a lower risk at the endpoint for the patient with the shorter stay. Including this loss function allows the model to optimise for concordance. The concordance was the only evaluation metric used in the paper which might have lead to potentially biased and inflated results. Subsequently reproduced work has shown that indeed this model is not calibrated well and significantly lags across other metrics when compared to simpler models [7, 10, 23, 32].
A-B5 Logistic Hazard
The Logistic Hazard method is a submodular implementation of our deep learning model using the loss in (4) with an MLP that similarly parametrises the PMF of the survival times [28]. DySurv expands on this method to include it as a component in the framework with the variational autoencoder to jointly optimise for both tasks of reconstruction and latent space formation as well as survival estimation. A key difference between the log-likelihood loss used here (and in our model) and in DeepHit is that logistic hazards do not allow for survival past the maximum time horizon.
A-B6 CoxTime
CoxTime is a relative risk model that extends Cox regression beyond the proportional hazards and is the first of the continuous-time methods. The standard Cox regression model which we will not spend space introducing here consists of a baseline hazard term (defined cumulatively in the loss by a pre-selected estimator such as Breslow) and a relative risk term which is an exponential factor of the weighted linear combination of features. The basic model assumes constant proportionality between the patients’ hazards over time and is thus restrictive. In other words, the difference between survival likelihoods for a given time is proportional to the difference in feature or hazard values for patients. CoxTime goes around this assumption by parametrising the relative risk term as a function of time and not just the features, thus the non-proportional behaviour over time is modelled by allowing for time to be considered alongside the features [5].
A-B7 CoxCC
CoxCC (Case Control) is just a proportional implementation of the CoxTime model and is the closest to the standard Cox implementation where the minimisation of the partial log-likelihood is done with stochastic gradient descent by averaging over constrained risk sets for mini-batch learning instead of the entire dataset like in classical survival analysis [5].
A-B8 DeepSurv
DeepSurv is a deep learning model that directly minimises the negative partial log-likelihood as defined in the standard Cox model. It is one of the first deep learning implementations for survival analysis and the output of the model is the log-risk term of the Cox model which accounts for nonlinearity [6]. There is no indirect estimation of cumulative risk or survival through likelihood estimation like in the previous methods, thus DeepSurv similarly suffers under limitations of the Cox such as the proportionality assumption [33].
A-B9 PCHazard
The last continuous-time method we introduce and implement is PCHazard which assumes that the continuous-time hazard rate (instantaneous value of risk) is piecewise constant. The method relies on optimising for the likelihood contribution which mimics the MTLR approach albeit in the continuous-time setting with the hazards parametrised by a simple MLP. The piecewise constant causes the likelihood to behave like a Poisson likelihood [34]. Despite the method operating in continuous-time, the hazards are defined in time intervals which rely on discretisation steps from the observed continuous event times and censoring times, while we discretise the times to a predefined set of time flagposts [19].
A-C Model architecture
Following the LSTM unit, the remaining part of the encoder consists of an MLP module with 3 layers. The encoder and decoder are mirrored in their structure with the hidden neurons in the MLP layers consisting of 3 times feature length, then 5 times, then 3 times, before passing onto the last output layer. For the encoder, the last layer is used in approximating the mean and standard deviation of the latent Gaussian distribution. As for the decoder, the last output layer is used for reconstructing the input vector. The survival module is similar to the MLP components of the encoder and decoder. All components are jointly optimised through multi-loss optimisation. All of the components have been investigated with and without dropout included. The output of the survival module is 10 nodes softmaxed hence properly jointly distributed, each of which gives the probability that the patient has suffered the event (death) at that specific time interval.
Figure 5: Proposed DySurv model for dynamic risk prediction based on a combination of static and time-series EHR data
A-D Ablation studies
To empirically arrive at the optimal configuration for the architecture, we undertook ablation studies and grid search combinations for hyperparameters using a validation set presented in the tables below. The hyperparameters investigated for DySurv include:
  • Dropout: 0.5, 0.6, 0.7, 0.8
  • Number of hidden layers: 2, 3, 4, 5
  • Number of hidden neurons: 10, 25, 75, 100, current
  • Latent representation vector length: 10, 20, 30
  • Learning rate: 0.00001, 0.0001, 0.001, 0.01, 0.1
  • Batch size: 8, 16, 32, 64, 128, 256, 512
TABLE XI: Ablation study experiments with DySurv on MIMIC-IV using time-series and static components with the optimal hyperparameter combination. The top strategy was kept as the optimal one when adding lower strategies. For example, once the inclusion of dropout was found to contribute positively to the results, it was included in the model architecture, then the addition of batch normalisation was tested.
$C_{ind}^{\mathrm{td}}$ IBS IBLL
Dropout ($p$ = 0.5)
With 57.9 0.122 0.320
Without 53.1 0.145 0.389
Batch Normalisation
With 57.6 0.130 0.322
Without 57.9 0.122 0.320
Conditional term concatenation
With 57.9 0.122 0.320
Without 55.7 0.144 0.377
DySurv ($+$ time-series) 57.9 0.122 0.320

### v3

@@
 Abstract
-Survival analysis focuses on estimating time-to-event distributions which can help in dynamic risk prediction in healthcare.
-Extending beyond the classical Cox model, deep learning techniques have been developed which moved away from the constraining assumptions of proportional hazards.
-Traditional statistical models often only include static information where, in this work, we propose a novel conditional variational autoencoder-based method called DySurv, which uses a combination of static and time-series measurements from patient electronic health records to estimate the risk of death dynamically.
-DySurv has been tested on several time-to-event benchmarks where it outperforms existing methods, including deep learning methods, and we evaluate it on real-world intensive care unit data from MIMIC-IV and eICU.
-The predictive capacity of DySurv is consistent and the survival estimates remain disentangled across different datasets supporting the idea that dynamic deep learning models based on conditional variational inference in multi-task cases can be robust models for survival analysis.
+Objective: Machine learning applications for longitudinal electronic health records often forecast the risk of events at fixed time points whereas survival analysis achieves dynamic risk prediction by estimating time-to-event distributions.
+Here, we propose a novel conditional variational autoencoder-based method, DySurv, which uses a combination of static and longitudinal measurements from electronic health records to estimate the individual risk of death dynamically.
+Materials and Methods: DySurv directly estimates the cumulative risk incidence function without making any parametric assumptions on the underlying stochastic process of the time-to-event.
+We evaluate DySurv on 6 time-to-event benchmark datasets in healthcare, as well as two real-world ICU EHR datasets extracted from eICU and MIMIC-IV.
+Results: DySurv outperforms other existing statistical and deep learning approaches to time-to-event analysis across concordance and other metrics.
+It achieves time-dependent concordance of over 60% in the eICU case.
+It is also over 12% more accurate and 22% more sensitive than in-use ICU scores like APACHE and SOFA.
+The predictive capacity of DySurv is consistent and the survival estimates remain disentangled across different datasets.
+Discussion: Our interdisciplinary framework successfully incorporates deep learning, survival analysis, and intensive care to create a novel method for time-to-event prediction from longitudinal health records.
+We test our method on several held-out test sets from a variety of healthcare datasets and compare it to existing in-use clinical risk scoring benchmarks.
+Conclusion: While our method leverages non-parametric extensions to deep learning-guided estimations of the survival distribution, further deep learning paradigms could be explored.
 Index Terms:
 deep learning, healthcare, personalized medicine, prognostication, survival analysis, variational autoencoders
-I Introduction
+I BACKGROUND
 Survival analysis refers to statistical approaches to estimating distributions of event times or times it takes for an event to happen as well as rates of survival over time while accounting for censoring.
@@
 However, the limitations of standard statistical models such as the Cox proportional hazards model include lack of complexity, constraining assumptions about the behaviour or proportions of covariate effects over time, and the reliance on only using static covariates [4].
-While there are extensions of Cox to time-varying covariates, these still suffer from the proportionality assumption (often violated in practice) and are often outperformed by simple deep learning models [5].
-In settings like healthcare and with the occurrence of electronic health records, high-dimensional longitudinal information on the patient’s state can be an informative source for the prediction of clinical risk of mortality and other events.
+While there are extensions of Cox to time-varying covariates, these still suffer from the proportionality assumption (often violated in practice) and are outperformed by simple deep learning models [5].
+The proportional hazards assumption inherent in Cox models states that the hazard rates of different patients over time remain in fixed proportion to each other, in other words, despite the potential change in risk trajectories for a patient in the future, these changes remain fixed relative to other patients.
+Thus, implying that the hazard in measuring the effect of any predictor is constant over time.
+This severely restraints the predicted risk trajectories and they often do not capture real changes in risk over time for a patient [6].
+In settings like healthcare, high-dimensional longitudinal information on the patient’s state can be an informative source for the prediction of clinical risk of mortality and other events.
 To incorporate longitudinal, high-dimensional, and potentially multimodal data and to move away from constraining assumptions of simple statistical models, deep learning methods have been proposed in survival analysis.
-These include extensions to the Cox model which estimate the parameters using a deep learning model that thus remain restrained with the proportionality assumption as in the case of DeepSurv [6].
+These include extensions to the Cox model which estimate the parameters using a deep learning model.
+However, these models remain restrained with the proportionality assumption as Cox, such as the case of DeepSurv [7].
+Existing methods often rely on parametric or semi-parametric assumptions of the survival distribution, thereby potentially restraining the predicted distribution away from a closer representation of the true survival distribution [8].
 Other deep learning models include DeepHit and its extension to longitudinal data, Dynamic-DeepHit, which implements a custom loss and a simple recurrent neural network to avoid making any parametric assumptions on the survival distribution [9, 10].
@@
 Our model, DySurv, is compatible with both static and longitudinal time-series data enabling more comprehensive learning from patient electronic health records.
-We validate our approach both in static and time-varying settings using six benchmark datasets as well as two large public open-access ICU datasets.
-The critical care setting provides high-frequency longitudinal measurements and short-term health outcomes in which dynamic risk stratification of survival analysis could provide additional insights [13].
+We validate our approach both in static and time-varying settings using six benchmark datasets as well as two large public open-access ICU datasets.The critical care setting provides high-frequency longitudinal measurements and short-term health outcomes in which dynamic risk stratification can help with urgent prognostication and prevention over just detection [15].
 We aim to show that conditional variational inference and autoencoder reconstruction tasks can improve learning from complex time-series data by extracting latent features to optimise the survival task.
+DySurv accomplishes this by adding a conditional variational inference loss to the cumulative risk loss estimation from logistic hazards which aids in learning the survival task.
 Since DySurv is not only multimodal in input but also in output with both reconstruction and survival tasks included, conditional variational autoencoders (CVAE) provide better predictive performance [16].
-By mapping high-dimensional covariates into a lower-dimensional space, CVAEs can effectively capture the underlying structure and identify hidden factors that significantly impact survival outcomes to predict mortality as well as time-to-death.
+By mapping high-dimensional covariates into a lower-dimensional space, CVAEs can effectively capture the underlying latent structure of the data to estimate survival outcomes.
 This capability allows deeper insights into dynamic mortality risk prediction in healthcare.
-II Materials and Methods
+II METHODS
 II-A Data
@@
 Quantile transformations have been applied for standardisation and fit only on the training dataset.
-Please see the appendix for a full list of features for each dataset.
+Please see the Supplementary for a full list of features for each dataset.
 II-A1 SUPPORT
@@
 The dataset consists of 8,873 samples and 14 static features.
+A total of 68.1 percent of patients died with a median death time of 58 days.
 II-A2 METABRIC
 The Molecular Taxonomy of Breast Cancer International Consortium contains genetic and clinical data from breast cancer patients with 1,904 samples and 9 static features [18].
+57.7% have an observed death with a median survival time of 116 months.
 II-A3 GBSG
-The Rotterdam & German Breast Cancer Study Group contains treatment and clinical data on 2,232 breast cancer patients with 6 static features [17].
+The Rotterdam & German Breast Cancer Study Group contains treatment and clinical data on 2,232 breast cancer patients with 6 static features [19]. 79.6% of patients have an observed event.
 II-A4 NWTCO
-The National Wilm’s Tumor dataset contains staging and clinical data on 4,028 Wilms’ tumour patients with 6 static features [18].
+The National Wilm’s Tumor dataset contains staging and clinical data on 4,028 Wilms’ tumour patients with 6 static features [20]. 16.0% of patients have an observed event.
 II-A5 sac3
-The simulated dataset contains discrete event times with 44 static features and 100,000 samples [19].
+The simulated dataset contains discrete event times with 44 static features and 100,000 samples [21]. 37% of samples have been censored.
 II-A6 sac_admin5
-The simulated dataset contains discrete event times with 5 static features and 50,000 samples [20].
+The simulated dataset contains discrete event times with 5 static features and 50,000 samples [22]. 37% of samples have been censored.
 II-A7 MIMIC IV and eICU
@@
 The dataset contains 71,935 samples of ICU stays with 33 static features (categorical features were one-hot encoded) and 65 time-varying features.
-Furthermore, we tested survival analysis models in the eICU Collaboration Research Database [pollard2018eicu].
+Furthermore, we tested survival analysis models in the eICU Collaboration Research Database [24].
 The eICU database was processed using postgreSQL and the pandas package.
@@
 Static variables include age, sex, admission unit, and others that did not have missingness.
-For the time-series variables, we use forward filling as clinicians in practice would only consider the last recorded measurement.
+For the time-series variables, we use forward filling as clinicians in practice would only consider the last recorded measurement and as has been done by previous work on these datasets [25, 26].
 If the first set of measurements is missing for some time-varying features, instead of dropping those features or patients, we backward fill from the closest measurement in the future.
-The processing of MIMIC-IV follows from previously published work on this dataset but is adapted for the survival scenario with the duration of stay in the ICU or the maximum time horizon for the event times defined as 10 days and time-series features taken in 72-hour timesteps [22].
-Each node of the neural network model then corresponds to a 1-day risk prediction in the ICU as the maximum time horizon is 10 days.
+10.0% of patients had an observed event of death.
+The pre-processing of MIMIC-IV and eICU follows from previously published work on these datasets but is adapted for the survival scenario with the duration of stay in the ICU or the maximum time horizon for the event times defined as 10 days and time-series features taken in 72-hour timesteps after resampling to 1-hour intervals [27].
+9.5% of patients had an observed event of death in this cohort.
 Censoring, as always, is defined as loss-to-follow-up or discharge.
@@
 where $f(t)$ is the probability density function of event time and
-$f(t)=\lim_{\Delta t\rightarrow 0}\frac{\operatorname{P}(t<T\leq t+\Delta t)}{\Delta t}$
+${f(t)=-\frac{dF}{dt}}$ (2)
 and $F(t)$ corresponds to the cumulative risk or incidence function $F(t)=\operatorname{P}(T\leq t)$.
@@
 A common assumption is that this censoring is not important and independent of the outcome of the study itself [29].
-Once we have an estimate of the risk, we can then simply obtain the survival function and curves by taking its negation from one as shown in equation (3).
+Once we have an estimate of the risk, we can then simply obtain the survival function and curves by subtracting the cumulative risk from 1 as shown in equation 1.
 Data for survival analysis contain three main sets of variables, the first is the feature set which can consist of static or time-series features (the latter having measurements at potentially different sampling frequencies), the time-to-event for the events in question or censoring, respectively, and the outcome label for an event [30].
@@
 $\mathcal{D}=\left\{\left(\mathcal{X}^{i},\mathcal{T}^{i},y^{i}\right)\right\}_{i=1}^{N}$ (3)
-with $\mathcal{X}^{i}$ representing the feature matrix of patient $i$, $\mathcal{t}$ being the time-to-event ie.
-the minimum of the event and censoring time, $y$ being the label for the outcome, and $N$ samples included.
-A feature matrix (in the time-series case) for patient $i$, $\mathcal{X}^{i}$, can be seen as
-$\mathcal{X}^{i}=\left\{\boldsymbol{x}_{1}^{i},\boldsymbol{x}_{2}^{i},\ldots,\boldsymbol{x}_{j}^{i}\right\}$ (3)
+with $\mathcal{X}^{i}$ representing the feature matrix of patient $i$, $\mathcal{T}$ being the time-to-event i.e., the minimum of the event and censoring time, $y\in\{\varnothing,1\}$ being the label for the outcome with $\varnothing$ representing right-censoring, and $N$ samples included.
+Static features have been expanded into time-series by replication padding matching the respective time-series component.
+Effectively, the feature matrix then consists only of time-series features.
+A feature matrix for patient $i$, $\mathcal{X}^{i}$, can be seen as
+$\mathcal{X}^{i}=\left\{\bm{x}_{1}^{i},\bm{x}_{2}^{i},\ldots,\bm{x}_{j}^{i}\right\}$ (4)
 where $j$ is the length of the time-series for $1\leq j<J^{i}$ where $J$ is the maximum time step with timestamps of measurements $\left[t_{1}^{i},t_{2}^{i},\ldots,t_{j}^{i}\right]$.
-$\boldsymbol{x}_{j}^{i}$ contains $M$ set of features $\left[x_{j,1}^{i},x_{j,2}^{i},\ldots,x_{j,M}^{i}\right]^{\top}$ for timestamp $t_{j}^{i}$.
+$\bm{x}_{j}^{i}$ contains $M$ set of features $\left[x_{j,1}^{i},x_{j,2}^{i},\ldots,x_{j,M}^{i}\right]^{\top}$ for timestamp $t_{j}^{i}$.
 We undertake estimation of the underlying cumulative risk by mainly optimizing the negative logarithmic likelihood of the joint distribution of the event time and outcome with right-censoring.
@@
 For censored patients, we capture the censoring time conditioned on the measurements recorded prior to the censoring.
-If we assume $\hat{a}_{t}=P\left(T=t\mid\mathcal{X}\right)$ (the output of the last layer node of the neural network module) represents the estimated probability of experiencing an event at time $t$, then the loss can be represented as
+If we assume $\hat{a}_{t}=P\left(T=t\mid\mathcal{X}\right)$ represents the estimated probability of experiencing an event at time $t$, then the loss can be represented as
 $\mathcal{L}_{1}=-$ $\sum_{i=1}^{N}\left[\mathbb{1}\left(y^{i}\neq\varnothing\right)\cdot\log\left(\frac{\hat{a}_{t^{i}}^{i}}{1-\sum_{y\neq\varnothing}\sum_{n\leq t_{Ji}^{i}}\hat{a}_{n}^{i}}\right)\right.$ (5)
@@
 where $\mathbb{1}$ is the indicator function.
-The first term represents optimising for uncensored patients and the second for those censored at the last noted time ie.
+The first term represents optimising for cumulative incidence risk accounting for uncensored patients and the second for those censored at the last noted time ie.
 alive.
 The estimated cumulative incidence risk, $\hat{F}$, is the probability that the patient dies at time $t$ or before, conditioned on all previous longitudinal measurements.
-By optimising for this loss, we estimate the actual risk distribution for each patient and a prediction can be made for arbitrary times across event times as outputs of the last layer.
-As data are complex (i.e., static and time-varying), we seek to learn a latent representation using a conditional variational autoencoder that would improve learning for the task of survival analysis.
-There are two loss terms, namely the reconstruction (mean squared error) and the Kullback–Leibler (KL) divergence in the variational autoencoder.
+By optimising for this loss, we estimate the actual risk distribution for each patient and a prediction can be made for arbitrary times across event times.
+As data are complex (i.e., static and time-varying), we seek to learn a latent representation using a conditional variational autoencoder that would improve learning for the task of survival analysis [31] [^1] For those readers less familiar with the autoencoder architecture and machine learning background for survival analysis, we recommend the following insightful surveys and chapters to peruse: [8, 32, 33].
+Since we are working with a deep learning model, our goal is to minimise a specific loss function relevant to our task.
+The loss will penalise our model when it starts learning away from the task in question.
+The first loss term relates to learning an efficient cumulative risk distribution to represent survival as described above in Equation 5.
+There are two further loss terms, namely the reconstruction (mean squared error) and the Kullback–Leibler (KL) divergence in the variational autoencoder [34].
+The first helps the model learn better representations of the input data, while the second helps learn better latent representations.
 Estimating the distribution of the underlying latent factors relies on minimising the KL divergence between an approximation of the true posterior and the true distribution both of which are assumed to be multivariate Gaussians.
@@
 $\alpha$ is considered as a hyperparameter that is optimised during training according to multi-objective optimization principles.
-Figure 5 in the Appendix shows the structure of the proposed model.
 DySurv leverages the dynamic nature of deep learning time-series models combined with conditional variational autoencoders for multi-task learning of survival analysis risk prediction extending beyond the classical fixed binary event prediction from traditional machine learning models.
@@
 Using a simple autoencoder has been shown to lead to overfitting and imbalanced learning of the reconstruction task that could harm learning the survival task whereas a VAE’s objective function is based on the reconstruction loss from a randomly sampled vector allowing for more robustness [14].
-We concatenate the feature vectors from the static and time-series features together before feeding them into an encoder equipped with a Long-Short Term Memory (LSTM) cell.
+We concatenate the feature vectors from the static and time-series features together before feeding them into an encoder equipped with a Long-short-term-memory (LSTM) cell.
+The LSTM module allows us to learn the temporal patterns in the longitudinal data and represent it in a hidden embedding for downstream tasks such as the estimation of the survival function.
+A short primer on the LSTM cell can be found in Supplementary Section C.
 The time-series components are compressed into a latent representation that is then used, as is, as input to the estimation of the negative log-likelihood loss function.
-A detailed description of the architecture can be found in the Appendix.
+A detailed description of the architecture can be found in the Supplementary.
 Once the latent vector is sampled from the Gaussian distribution defined by these parameters, the lower-dimensional latent factors are used as input for an MLP module to optimize the survival task.
+In the case of the ICU datasets, for example, each output node of this MLP, a representation of $\hat{a}_{t}$, then corresponds to a 1-day risk prediction in the ICU for a maximum time horizon of 10 days.
 The hyperparameters optimised in the network through grid search using the training and validation set included learning rate, batch size, $\alpha$, and dropout proportion.
-To minimise overfitting, we employ early stopping techniques.
+To minimise overfitting, we employ early stopping.
 Figure 1: (a) Description of the proposed DySurv framework using longitudinal EHR data for dynamic risk prediction instead of fixed-point event classification.
@@
 at all reasonable times into the feature.
-We compare the performance of our model to a collection of survival analysis models summarised in Table I and described in detail in the Appendix.
+We compare the performance of our model to a collection of survival analysis models summarised in Table I and described in detail in the Supplementary.
+Unlike DySurv, existing methods like Cox-based methods as well as some deep learning alternatives, rely on assuming a specific statistical distribution of the survival distribution, thus making parametric distributions that restrict their predictive power.
+This assumption is described in further detail in Supplementary Section B.
 Table I: Survival analysis methods investigated
@@
 Hence, we will rely on using the time-dependent extension $C_{ind}^{\mathrm{td}}$ with some modifications accounting for predictions independent of feature observations having a concordance of 0.5.
+The metric was originally proposed in [38] and all of the included metrics are provided in the PyCox package.
 The metric can be represented as
-$C_{ind}^{\mathrm{td}}=\mathrm{P}\left\{\hat{S}\left(T_{i}\mid\mathbf{x}_{i}\right)<\hat{S}\left(T_{i}\mid\mathbf{x}_{j}\right)\mid T_{i}<T_{j},D_{i}=1\right\}$ (8)
-where $\hat{S}$ indicates the estimated survival probabilities are used and $D_{i}=1$ that only those who experienced the event are considered in this metric.
+$C_{ind}^{\mathrm{td}}=\mathrm{P}\left\{\hat{S}\left(T^{i}\mid\mathbf{x}^{i}\right)<\hat{S}\left(T^{i}\mid\mathbf{x}^{j}\right)\mid T^{i}<T^{j},y^{i}=1\right\}$ (9)
+where $\hat{S}$ indicates the estimated survival probabilities are used and $y^{i}=1$ that only those who experienced the event are considered in this metric.
 A noted limitation of this metric is its obvious bias and dependence on the censoring distribution as only non-censored samples are considered making it affected by the length of stay and the censoring proportion that increases over the length of stay.
@@
 The equation used is thus
-$\operatorname{BS}_{\operatorname{IPCW}}(t)=\frac{1}{n}\sum_{i=1}^{n}[\frac{\hat{S}_{i}(t)^{2}1\left\{T_{i}\leq t,D_{i}=1\right\}}{\hat{G}_{i}\left(T_{i}-\right)}+$ (9)
-$\frac{\left[1-\hat{S}_{i}(t)\right]^{2}1\left\{T_{i}>t\right\}}{\hat{G}_{i}(t)}]$
-where $\hat{G}_{i}(t)=\mathrm{P}\left(C_{i}^{*}>t\right)>0$ is the Kaplan-Meier estimate of the censoring distribution for sample $i$ and $C_{i}^{*}$ is the censoring time.
+$\operatorname{BS}_{\operatorname{IPCW}}(t)=\frac{1}{n}\sum_{i=1}^{n}[\frac{\hat{S}^{i}(t)^{2}\mathbb{1}\left\{T^{i}\leq t,y^{i}=1\right\}}{\hat{G}^{i}\left(T^{i}-\right)}+$ (10)
+$\frac{\left[1-\hat{S}^{i}(t)\right]^{2}\mathbb{1}\left\{T^{i}>t\right\}}{\hat{G}^{i}(t)}]$
+where $\hat{G}^{i}(t)=\mathrm{P}\left(C^{i*}>t\right)>0$ is the Kaplan-Meier estimate of the censoring distribution for sample $i$ and $C^{i*}$ is the censoring time.
 The expected value of this metric is the same as that for the uncensored Brier Score.
@@
 A limitation of this metric, however, is the biased assumption of the censoring distribution being the same across samples thereby disregarding covariate effects.
-This can be addressed by using an administrative extension of the metric that requires access to all the censoring times but a discussion of this in greater detail can be perused here [20].
-Lastly, we introduce the IPCW (negative) binomial log-likelihood or NBLL from classic binary classification which measures both discrimination and calibration of the estimates and uses its integrated extension for all times INBLL
-$\operatorname{BLL}(t)=\frac{1}{N}\sum_{i=1}^{N}[\frac{\log\left[1-\hat{S}\left(t\mid\mathbf{x}_{i}\right)\right]1\left\{T_{i}\leq t,D_{i}=1\right\}}{\hat{G}\left(T_{i}\right)}+$ (11)
-$\frac{\log\left[\hat{S}\left(t\mid\mathbf{x}_{i}\right)\right]1\left\{T_{i}>t\right\}}{\hat{G}(t)}]$
+This can be addressed by using an administrative extension of the metric that requires access to all the censoring times but a discussion of this in greater detail can be perused where the metric was originally proposed [22].
+Lastly, we introduce the IPCW (negative) binomial log-likelihood or NBLL from classic binary classification from [22] which measures both discrimination and calibration of the estimates and uses its integrated extension, INBLL, for all times
+$\operatorname{BLL}(t)=\frac{1}{N}\sum_{i=1}^{N}[\frac{\log\left[1-\hat{S}\left(t\mid\mathbf{x}^{i}\right)\right]1\left\{T^{i}\leq t,y^{i}=1\right\}}{\hat{G}\left(T^{i}\right)}+$ (12)
+$\frac{\log\left[\hat{S}\left(t\mid\mathbf{x}^{i}\right)\right]1\left\{T^{i}>t\right\}}{\hat{G}(t)}]$
 […]
@@
 We provide survival curves for a random group of five samples/patients (with different event times) across all datasets that show a clear separation of risk as can be seen in Figure 2.
-For MIMIC-IV, the data input consists of both static and time-series data and we provide survival curve results for both scenarios as well as for Dynamic-DeepHit for comparison.
-DySurv has a better capacity to separate independent patient risk trajectories as shown in Figure 3.
-(a) Survival curve (estimate of survival probability over time) for five random samples from SUPPORT
-(b) Survival curve (estimate of survival probability over time) for five random samples from METABRIC
-(c) Survival curve (estimate of survival probability over time) for five random samples from GBSG
-(d) Survival curve (estimate of survival probability over time) for five random samples from NWTCO
-(e) Survival curve (estimate of survival probability over time) for five random samples from sac3
-(f) Survival curve (estimate of survival probability over time) for five random samples from sac5
-Figure 2: The set of survival curves for benchmark datasets as generated by DySurv shows clear extrapolation of risk across different samples.
-Since DySurv provides discrete estimates over time-intervals, the curves were generated with additional interpolation commonly used in other discrete methods in survival analysis.
+For MIMIC-IV, the data input consists of both static and time-series data and we provide survival curve results for both scenarios in Figure 3.
+(a) Survival curve for five random samples from SUPPORT
+(b) Survival curve for five random samples from METABRIC
+(c) Survival curve for five random samples from GBSG
+(d) Survival curve for five random samples from NWTCO
+(e) Survival curve for five random samples from sac3
+(f) Survival curve for five random samples from sac5
+Figure 2: Survival curves (estimate of survival probability over time) for benchmark datasets by DySurv across different samples.
+DySurv provides discrete estimates over time and additional interpolation was applied.
+Dots correspond to true event times which were predicted correctly by DySurv.
 (a) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using only static features with DySurv
 (b) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using both static and time-series features with DySurv
-(c) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using both static and time-series features with Dynamic-DeepHit
-Figure 3: The set of survival curves for the MIMIC-IV ICU EHR dataset as generated by DySurv that shows clear extrapolation of risk across different patients as compared to both static, time-series, and Dynamic-DeepHit cases.
-The same set of patients was used in generating the survival curves across cases.
+Figure 3: The set of survival curves for the MIMIC-IV ICU EHR dataset as generated by DySurv shows extrapolation of risk across different patients as compared to static and time-series feature sets.
+Dots correspond to true event times which were predicted correctly by DySurv.
 Upon deployment of a trained model, DySurv can generate risk estimates through time for each patient while using their history of observations.
@@
 We also compare DySurv for 24-hour risk prediction to existing ICU survival scores in practice like APACHE IV and SOFA scores which can be seen in Figure 4.
-Figure 4: Comparison of DySurv at 24-hour prediction with existing ICU survival scores using (a) AUROC and (b) Sensitivity.
+Figure 4: Comparison of DySurv at 24-hour prediction with existing ICU survival scores and deep learning survival models on MIMIC-IV using (a) AUROC and (b) Sensitivity and c) AUPRC.
 APACHE IV was only able to be retrieved from MIMIC-IV and SOFA score was calculated from the eICU dataset.
-The calculations are available in the code repository.
-IV Discussion
+IV CONCLUSIONS
 The first set of results from Table II relate to applying DySurv only on static data from several benchmark datasets of varying sizes.
@@
 As for MIMIC-IV and eICU results, we see that even with only static data in the input, DySurv manages to outperform other survival analysis models.
-When time-series data is added to the input in a multi-modal fashion, it can help the model improve its performance and outperform both CoxTime and Dynamic-DeepHit across all metrics.
-The survival curves in Figure 3 show clearly that the static version of DySurv has the clearest separation of survival trajectories for different patients and that the addition of time-series data makes the task more complex.
-This is probably due to the reconstruction task being a lot more difficult now that time-series are involved and the model makes sacrifices on the survival task front.
-However, compared to Dynamic-DeepHit, whose survival curves in Figure 3 c) are inseparable, the time-series version of DySurv still provides a relatively reliable survival estimate.
-We also see that, as we would expect clinically, the survival of the patients significantly changes in the last few days in the ICU, starting to drop a few days before death.
-Previous work has shown that earlier times in the ICU correspond to higher survival rates.
-This suggests that identifying the time period when survival rates drop dramatically in the ICU can help target earlier treatment for those most at risk [31].
-V Conclusion
+When time-series data are added to the input in a multi-modal fashion, it can help the model improve its performance and outperform both CoxTime and Dynamic-DeepHit across all metrics.
+Figure 3 shows examples of projected survival trajectories using DySurv in both the static-only and static-and-time-series feature settings.
+We also see that, as expected in the hospitalised setting, the survival of the patients significantly changes in the last few days in the ICU, starting to drop a few days before death.
+Previous work has shown that earlier discharge times in the ICU correspond to higher survival rates.
+This suggests that identifying the period when survival rates drop dramatically in the ICU can help target earlier treatment for those most at risk [41].
+We assume that patients will stay in the ICU for a maximum of 10 days and while that is generally true, there could be outliers and differences in ICU stay duration distributions.
+We suspect this would not severely hinder the generalizability of DySurv to those scenarios as that would just mean an expansion of the time component for the survival trajectory predictions.
+From Figure 4 we can see how DySurv outperforms existing in-use systems for patient risk prognosis in the ICU across a variety of metrics.
+We acknowledge the limitation of the results for generalisability from using different data subsets.
+The model is highly dependent on the quality and completeness of time-series data.
+While forward and backward filling has been commonly employed and verified with clinical collaborators [26, 25], this approach could introduce biases, particularly if the missingness is non-random or systematically tied to patient outcomes.
+Furthermore, we used pre-defined time intervals (72-hour windows) which may not fully capture the dynamic variability in patient trajectories, particularly for patients with more erratic health progressions.
+This time interval, however, can be manually adjusted depending on the identified application.
+We also acknowledge potential overfitting to the benchmark datasets, especially given the relatively small size of some of these datasets.
+This could hinder generalization to more diverse populations or clinical settings where patient characteristics and ICU protocols differ significantly.
+We have found that careful hyperparameter optimisation can aid in addressing this challenge when learning on different datasets.
+The generalizability of DySurv is also constrained by the assumptions embedded in the model architecture, despite being nonparametric.
+While DySurv does not assume proportional hazards, it still relies on the assumptions inherent in its logistic hazards model producing discrete estimates of the survival probabilities.
+Additionally, as deep learning models are inherently opaque, their extensions to healthcare settings come with concerns of trustworthiness and interpretability.
+To address this limitation and further verify that DySurv is learning relevant clinical patterns from the patient data, we used permutation importance to obtain the top 10 predictive features for DySurv on MIMIC-IV and eICU.
+These results are in line with relevant clinical literature and practice and a discussion can be found in Section G in the Supplementary.
 In this paper, we present a novel dynamic risk prediction model for survival analysis based on deep learning in survival analysis paradigms.
+We work collaboratively with intensive care physicians and demonstrate interdisciplinary pathways to the implementation of deep learning for survival analysis methods in the ICU as shown in Figure 2 in the Supplementary.
 Our method builds on a combination of previous work including Dynamic-DeepHit by leveraging direct learning of the joint distribution of the first event time and the event through log-likelihood optimisation with logistic hazards.
-Theoretically, this approach is an alternative to the risk log-likelihood loss function of Dynamic-DeepHit itself that does not use a ranking loss for biased inflation of concordance results.
+Theoretically, this approach is an alternative to the risk log-likelihood loss function of Dynamic-DeepHit itself, which does not use a ranking loss for biased inflation of concordance results.
 Our DySurv model is capable of learning from complex EHR ICU time-series data and extracting lower-dimensional latent representations that can be useful for learning the survival task while also balancing reconstruction.
-As the model has been difficult to train due to loss instabilities and sensitivity to hyperparameter selection, future work can explore including a regularisation component to the loss terms.
+As the model has been difficult to train due to loss instabilities and sensitivity to hyperparameter selection, future work can explore including a regularization component to the loss terms.
 By using the underlying latent distribution, we can directly model an alternative to the survival distribution like the Weibull distribution instead of using a Gaussian intermediate.
-VI Conflicts of interest
-The authors declare no competing interests.
-VII Code and Data sharing
-The data used in the study are all publicly available.
-MIMIC-IV can be accessed after some initial requests and approvals from the PhysioNet website.
-The remaining datasets have their sources listed in the PyCox library repository.
-All of the code will be made available in a public GitHub repository after publication, including easily accessible notebooks to reproduce experiments from the paper.
-VIII Acknowledgements
-Munib Mesinovic appreciates the support of the Rhodes Trust and the EPSRC Centre for Doctoral Training in Health Data Science.
-T.Z.
-was supported by the Royal Academy of Engineering under the Research Fellowship scheme.
-Appendix A Appendix
-A-A Data Description
-A detailed list of features used in the study from different datasets can be seen in Tables V, VI, VII, VIII and IX.
-TABLE V: Features extracted from the SUPPORT survival dataset.
-Feature Type Feature Type
-Sex binary Age integer
-Number of comorbidities integer SUPPORT Coma Score continuous
-Years of Education continuous APS III Score continuous
-Diabetes binary Dementia binary
-3rd Day MAP continuous 3rd Day WBC continuous
-3rd Day HR continuous 3rd Day RR continuous
-Temperature continuous 3rd Day Bilirubin continuous
-TABLE VI: Features extracted from the NWTCO survival dataset.
-Feature Type Feature Type
-Relapse binary Age integer
-Histology Local binary Histology Institute binary
-Tumor Stage categorical Study Type categorical
-TABLE VII: Features extracted from the GBSG survival dataset.
-Feature Type Feature Type
-Hormone Treatment binary Age integer
-Tumor Grade categorical Menopause Treatment binary
-Progesterone (fmol) continuous Number of positive nodes categorical
-Estrogen (fmol) continuous
-TABLE VIII: Features extracted from the METABRIC survival dataset.
-Feature Type Feature Type
-Age integer Gene I binary
-Gene II binary Gene III binary
-Hormone Treatment binary Gene IV binary
-Radiotherapy binary Chemotherapy binary
-ER-positive binary
-TABLE IX: Features extracted from the MIMIC-IV database.
-The features include demographic data collected for all patients, ICU unit-specific information like the type of unit, hospital information like regional location, time since admission, vital signs including respiratory rate and blood pressure, and biochemical measurements including blood glucose and haemoglobin.
-Static Variables
-Feature Type Feature Type
-Sex binary Motor Response continuous
-Age integer Verbal Response continuous
-Height continuous Ethnicity categorical
-Weight continuous Unit Type categorical
-Hour of Admission integer Admission Location categorical
-Time Since Admission continuous
-Time-series Variables
-Feature Type Feature Type
-GCS - Eye continuous
-Eye Response continuous GCS - Motor continuous
-Braden Score continuous GCS - Verbal continuous
-Strength L Arm continuous Daily Weight continuous
-Strength R Arm continuous ALT continuous
-Strength L Leg continuous AST continuous
-Strength R Leg continuous HCO3 continuous
-Insurance categorical Hct continuous
-ALT continuous Alkaline Phosphatase continuous
-Anion Gap continuous AST continuous
-Base Excess continuous Bicarbonate continuous
-Bilirubin continuous Calcium continuous
-Total CO2 continuous Chloride continuous
-Creatinine continuous Glucose continuous
-Hematocrit continuous Hemoglobin continuous
-INR(PT) continuous Lactate continuous
-MCH continuous MCHC continuous
-MCV continuous Magnesium continuous
-PT continuous PTT continuous
-Phosphate continuous Platelet Count continuous
-Potassium continuous RDW continuous
-Red Blood Cells continuous Sodium continuous
-Urea Nitrogen continuous White Blood Cells continuous
-pCO2 continuous pH continuous
-pO2 continuous JH-HLM continuous
-Dyspnea Assessment continuous Daily Weight continuous
-Glucose continuous Heart Rate continuous
-DBP continuous SBP continuous
-O2 Flow continuous O2 Sat (%) continuous
-Pain Level continuous Pain Level Response continuous
-Phosphorous continuous Respiratory Rate continuous
-Richmond-RAS Scale continuous Temperature (${}^{\circ}$) continuous
-TABLE X: Features extracted from the eICU database.
-The features include demographic data collected for all patients, ICU unit-specific information like type and number of beds, hospital information like regional location and teaching status, vital signs including respiratory rate and blood pressure, and biochemical measurements including troponin and levels of potassium and protein in the blood.
-Feature Type Feature Type
-Sex binary Unit Stay Type categorical
-Age integer Num Beds Category categorical
-Height continuous Region categorical
-Weight continuous Teaching Status binary
-Ethnicity categorical Physician Speciality categorical
-Unit Type categorical Unit Type categorical
-Unit Admit Source categorical Mechanical Ventilation binary
-Unit Visit Number categorical
-Time-series Variables
-Feature Type Feature Type
-Base Excess continuous
--basos continuous FiO2 continuous
--eos continuous HCO3 continuous
--monos continuous Hct continuous
--polys continuous Hgb continuous
-ALT continuous MCH continuous
-AST continuous MCHC continuous
-BUN continuous MCV continuous
-O2 Sat (%) continuous MPV continuous
-PT-INR continuous PT continuous
-RBC continuous PTT continuous
-RDW continuous WBC continuous
-Alkaline ph. continuous Albumin continuous
-Bedside Glucose continuous Anion Gap continuous
-Calcium continuous Bicarbonate continuous
-Creatinine continuous Glucose continuous
-Lactate continuous Magnesium continuous
-pH continuous paCO2 continuous
-paO2 continuous Phosphate continuous
-Platelets continuous Potassium continuous
-Sodium continuous Bilirubin continuous
-Protein continuous Troponin - I continuous
-Urinary s. Gravity continuous mean BP continuous
-SBP continuous DBP continuous
-A-B Benchmark Models
-To compare DySurv to other survival analysis methods, we implement a selection of the most popular and consistently cited methods in the field and provide a short description of each below.
-We will first introduce discrete-time methods which rely on discretising the event times into specified durations, and then follow with continuous-time methods.
-A-B1 PMF
-The parametrisation of the Probability Mass Function (PMF) of the event times is another way of estimation without resorting to using discrete-time risk or hazard in likelihood optimisation.
-We described the continuous probability density function without using discrete time boundaries.
-It is the foundation of other methods like DeepHit and Multi-Task Logistic Regression.
-It similarly resorts to optimising a negative log-likelihood loss but instead of using the cumulative risk function, it uses the approximations of the PMF and survival functions [19].
-Since we can establish a direct representational relation between the risk and survival functions, the PMF loss can be seen as an alternative to our loss function.
-In our and others’ implementation, the PMF method is a simple Multi-Layer Perceptron (MLP) optimised for this loss and we use the same structure as we use for our survival module to be described later.
-A-B2 MTLR
-Multi-Task Logistic Regression provides a generalization of the binomial log-likelihood to jointly model the sequence of binary labels for each time interval risk prediction.
-This method similarly minimises the negative log-likelihood with the PMF and survival function terms but the network outputs are cumulatively summed in reverse to no certain advantage and, in fact, just add computational complexity and numeric instability [27, 29].
-A-B3 BCESurv
-This Binary Cross-Entropy for Survival is a method consisting of a set of binary classifiers that remove individuals as they are censored.
-The loss is the binary cross entropy of the survival estimates at a set of discrete times, with targets that are indicators of surviving each time.
-Each output node in the last layer corresponds to a binary classifier evaluated at that time point.
-As censored patients are removed, the method is biased towards those with higher event probabilities [20].
-A-B4 DeepHit
-The single-risk version of DeepHit is a deep learning model whose output nodes are softmaxed to jointly model the probabilities between the event times and the time durations are discretised like in our case.
-The model depends on optimising both the negative log-likelihood loss based on the cumulative incidence function and a ranking loss built on the intuition of the concordance.
-The ranking loss penalises incorrect ordering of patient pairs in which the patient that remains longer in the study should have a lower risk at the endpoint for the patient with the shorter stay.
-Including this loss function allows the model to optimise for concordance.
-The concordance was the only evaluation metric used in the paper which might have lead to potentially biased and inflated results.
-Subsequently reproduced work has shown that indeed this model is not calibrated well and significantly lags across other metrics when compared to simpler models [7, 10, 23, 32].
-A-B5 Logistic Hazard
-The Logistic Hazard method is a submodular implementation of our deep learning model using the loss in (4) with an MLP that similarly parametrises the PMF of the survival times [28].
-DySurv expands on this method to include it as a component in the framework with the variational autoencoder to jointly optimise for both tasks of reconstruction and latent space formation as well as survival estimation.
-A key difference between the log-likelihood loss used here (and in our model) and in DeepHit is that logistic hazards do not allow for survival past the maximum time horizon.
-A-B6 CoxTime
-CoxTime is a relative risk model that extends Cox regression beyond the proportional hazards and is the first of the continuous-time methods.
-The standard Cox regression model which we will not spend space introducing here consists of a baseline hazard term (defined cumulatively in the loss by a pre-selected estimator such as Breslow) and a relative risk term which is an exponential factor of the weighted linear combination of features.
-The basic model assumes constant proportionality between the patients’ hazards over time and is thus restrictive.
-In other words, the difference between survival likelihoods for a given time is proportional to the difference in feature or hazard values for patients.
-CoxTime goes around this assumption by parametrising the relative risk term as a function of time and not just the features, thus the non-proportional behaviour over time is modelled by allowing for time to be considered alongside the features [5].
-A-B7 CoxCC
-CoxCC (Case Control) is just a proportional implementation of the CoxTime model and is the closest to the standard Cox implementation where the minimisation of the partial log-likelihood is done with stochastic gradient descent by averaging over constrained risk sets for mini-batch learning instead of the entire dataset like in classical survival analysis [5].
-A-B8 DeepSurv
-DeepSurv is a deep learning model that directly minimises the negative partial log-likelihood as defined in the standard Cox model.
-It is one of the first deep learning implementations for survival analysis and the output of the model is the log-risk term of the Cox model which accounts for nonlinearity [6].
-There is no indirect estimation of cumulative risk or survival through likelihood estimation like in the previous methods, thus DeepSurv similarly suffers under limitations of the Cox such as the proportionality assumption [33].
-A-B9 PCHazard
-The last continuous-time method we introduce and implement is PCHazard which assumes that the continuous-time hazard rate (instantaneous value of risk) is piecewise constant.
-The method relies on optimising for the likelihood contribution which mimics the MTLR approach albeit in the continuous-time setting with the hazards parametrised by a simple MLP.
-The piecewise constant causes the likelihood to behave like a Poisson likelihood [34].
-Despite the method operating in continuous-time, the hazards are defined in time intervals which rely on discretisation steps from the observed continuous event times and censoring times, while we discretise the times to a predefined set of time flagposts [19].
-A-C Model architecture
-Following the LSTM unit, the remaining part of the encoder consists of an MLP module with 3 layers.
-The encoder and decoder are mirrored in their structure with the hidden neurons in the MLP layers consisting of 3 times feature length, then 5 times, then 3 times, before passing onto the last output layer.
-For the encoder, the last layer is used in approximating the mean and standard deviation of the latent Gaussian distribution.
-As for the decoder, the last output layer is used for reconstructing the input vector.
-The survival module is similar to the MLP components of the encoder and decoder.
-All components are jointly optimised through multi-loss optimisation.
-All of the components have been investigated with and without dropout included.
-The output of the survival module is 10 nodes softmaxed hence properly jointly distributed, each of which gives the probability that the patient has suffered the event (death) at that specific time interval.
-Figure 5: Proposed DySurv model for dynamic risk prediction based on a combination of static and time-series EHR data
-A-D Ablation studies
-To empirically arrive at the optimal configuration for the architecture, we undertook ablation studies and grid search combinations for hyperparameters using a validation set presented in the tables below.
-The hyperparameters investigated for DySurv include:
-  • Dropout: 0.5, 0.6, 0.7, 0.8
-  • Number of hidden layers: 2, 3, 4, 5
-  • Number of hidden neurons: 10, 25, 75, 100, current
-  • Latent representation vector length: 10, 20, 30
-  • Learning rate: 0.00001, 0.0001, 0.001, 0.01, 0.1
-  • Batch size: 8, 16, 32, 64, 128, 256, 512
-TABLE XI: Ablation study experiments with DySurv on MIMIC-IV using time-series and static components with the optimal hyperparameter combination.
-The top strategy was kept as the optimal one when adding lower strategies.
-For example, once the inclusion of dropout was found to contribute positively to the results, it was included in the model architecture, then the addition of batch normalisation was tested.
-$C_{ind}^{\mathrm{td}}$ IBS IBLL
-Dropout ($p$ = 0.5)
-With 57.9 0.122 0.320
-Without 53.1 0.145 0.389
-Batch Normalisation
-With 57.6 0.130 0.322
-Without 57.9 0.122 0.320
-Conditional term concatenation
-With 57.9 0.122 0.320
-Without 55.7 0.144 0.377
-DySurv ($+$ time-series) 57.9 0.122 0.320
+V Funding Statement
+This work was supported by the Rhodes Trust and the EPSRC Centre for Doctoral Training in Health Data Science grant (EP/S02428X/1).
+VI Competing Interests Statement
+The authors have no competing interests to declare.
+VII Contributorship Statement
+MM, PW, and TZ have all contributed to the conception or design of the work; or the acquisition, analysis, or interpretation of data for the work.
+MM has extracted and pre-processed the data with guidance from PW and TZ.
+MM has developed the method with guidance from TZ, designed the experiments, validated the results, and summarised the findings.
+MM, PW, and TZ drafted the work or reviewed it critically for important intellectual content.
+MM provided the initial draft and implemented the subsequent changes in writing.
+MM, PW, and TZ approved the final version for publication.
+All authors agree to be accountable for all aspects of the work in ensuring that questions related to the accuracy or integrity of any part of the work are appropriately investigated and resolved.
+VIII DATA AVAILABILITY
+The data is available from public access requests for eICU and MIMIC-IV.
+Accessing the data requires ethics module training and certification.
+It can be obtained at the following links: https://physionet.org/content/eicu-crd/2.0/ and https://physionet.org/content/mimiciv/2.0/.